tan [sin(4/5)]=x
Find x

Answer by Alan3354(69443)

tan [sin(4/5)]=x
Find x
--------------
r = 5
y = 4
--> x = 3 or -3
------
tan = y/x = 4/3 in Q1
= -4/3 in Q2